# ALPy Cogenesis

### v1

@@ V Conclusion
We have proposed a novel leptogenesis scenario by considering axion or ALP type interactions involving heavy and light chiral fermions such that the heavy fermion decay into the lighter counterpart and axion can generate a non-zero CP asymmetry. In the first realisation of this proposal, we consider two types of chiral fermions $N_{R},\nu_{R}$ such that out-of-equilibrium CP violating decay $N_{R}\rightarrow\nu_{R}a$ generates a net asymmetry in $\nu_{R}$ which later gets transferred to lepton doublets via Yukawa interactions mediated by a neutrinophilic Higgs doublet. Lepton number violation due to heavy $N_{R}$ eventually generates Majorana mass of $\nu_{R}$ at one loop level which results in type-I seesaw origin of light neutrino mass. Demanding this low scale type-I seesaw not to wash out $N_{R}$-generated asymmetry or generate a new asymmetry, we find the parameter space in terms of the lightest $N_{R}$’s mass $M_{1}$ and axion decay constant $f$. Given astrophysical bounds on $f\gtrsim 10^{8}$ GeV, we show that successful ALPy cogenesis can occur for $M_{1}\gtrsim 10$ TeV. For QCD axion, this keeps the detection prospects alive at different experiments searching for axions in $\mu$eV-eV range. Interestingly this range also accommodates axion DM. We also propose a Dirac version of the minimal ALPy cogenesis scenario which does not require any net lepton number violation. This not only requires the decaying heavy fermion to be of Dirac type but also require another light Dirac fermion $\chi$. Since asymmetric Dirac fermion $\chi$ can also contribute partially to dark matter in the Universe, the Dirac ALPy cogenesis has a wider parameter space to be probed by axion detection experiments.
Both the realisations proposed here can be probed at experiments operating at different frontiers. In addition to axion search experiments mentioned above, thermalised axion or light Dirac neutrinos can be probed via precision CMB measurements. Long-lived asymmetric dark matter, new sterile fermions as well as the neutrinophilic Higgs doublet can have rich phenomenology and detection prospects. While the phenomenology of flavor violating ALP couplings to the SM fermions have been studied in the literature MartinCamalich:2020dfe , the couplings introduced in ALPy cogenesis here can have additional UV completions and related phenomenology. We leave these details for future studies.

### v2

@@ V Conclusion
[…]
Both the realizations proposed here can be probed at experiments operating at different frontiers. In addition to axion search experiments mentioned above, thermalized axion or light Dirac neutrinos can be probed via precision CMB measurements. Long-lived asymmetric dark matter, new sterile fermions as well as the neutrinophilic Higgs doublet can have rich phenomenology and detection prospects. While the phenomenology of flavor violating ALP couplings to the SM fermions have been studied in the literature MartinCamalich:2020dfe , the couplings introduced in ALPy cogenesis here can have additional UV completions and related phenomenology. We leave these details for future studies.